The following units of a particular item were available for sale during the year:
Beginning inventory 150 units at $75
Sale 120 units at $125
First purchase 400 units at $78
Sale 200 units at $125
Second purchase 300 units at $80
Sale 290 units at $125
The firm uses the perpetual inventory system, and there are 240 units of the item on
hand at the end of the year. What is the total cost of the ending inventory according
to (a) FIFO, (b) LIFO?
1
Expert's answer
2014-11-14T14:42:33-0500
Beginning inventory 150 units at $75
Sale 120 units at $125
First purchase 400 units at $78
Sale 200 units at $125
Second purchase 300 units at $80
Sale 290 units at $125
Ending inventory 240 units.
What is the total cost of the ending inventory:
(a) According to FIFO:
Ending cost = 240*125 = $30,000
(b) According to LIFO:
Ending cost = 150*75 + 90*125 = $22,500